Errata for “Amarts: A class of asymptotic martingales; A. Discrete parameter” G. A. Edgar & Louis Sucheston

Page 197 Line 12. Replace σ ≥ τ_{0} by σ ≥ τ_{1} Page 197 Line 13. Replace [σ ≤ τ_{0}] by [σ ≤ τ_{1}] Page 197 Line 16. Replace σ_{1} ∈ T by σ_{1}, σ_{2} ∈ T Page 197 Line 17. Replace display by

σ_{1} = τ_{1} on A  σ_{2} = σ on A
= σ on A^{c}  = τ_{1} on A^{c}

Page 198 Line 3. Replace Y_{σ_{1}} by Y_{σ_{2}} Page 199 Line 16. Replace ℱ _{k} by (ℱ _{n})

Typeset by AMS-TEX